OpenStudy (trojanpoem):

Problem: y = tan(2arctanx) prove that: (1+x^2)y'' = 2(2y-x)y'

OpenStudy (trojanpoem):

I manged to get a solution for it, but I wanna know what's wrong on this solution.

OpenStudy (trojanpoem):

y = 2tan(arctanx)/1-tan^2(arctanx) = 2x/1 - x^2 (1-x^2)y = 2x -2xy + (1-x^2)y' = 2 -2y -2xy' + (-2x)y' + (1-x^2)y'' = 0 (1-x^2)y'' = 2(y - 2x)y'